# Three-Stream Temporal-Shift Attention Network Based on Self-Knowledge Distillation for Micro-Expression Recognition

### v2

@@ 4.3 Evaluation Metrics
For all experiments on the four public datasets, we employ leave-one-subject-out (LOSO) cross-validation. The unweighted accuracy (UAR) and unweighted F1 score (UF1) are used to evaluate the MER results. For convenience, $T{{P}_{i}}$, $F{{P}_{i}}$, and $F{{N}_{i}}$ are used to represent the predicted numbers of true positives, false positives, and false negatives for each class. $UF1$ and $UAR$ are calculated as follows:
@@
where $C$ represents the number of emotion categories in the dataset, and ${{N}_{i}}$ denotes the total number of MEs in the class ${{i}_{th}}$.
@@
STSTNet (Liong et al.,, 2019) 3 83.82 86.86
OFF-ApexNet (Gan et al.,, 2019) 3 86.97 88.28
$\mu$-BERT (Nguyen et al.,, 2023) 3 90.34 89.14
HTNet (Wang et al.,, 2023) 3 95.32 95.16
SKD-TSTSAN (ours) 3 99.13 98.67
@@ 4.4 Comparative Experiments Results
In this section, we compare the results of our method with other SOTA methods using four public ME datasets (CASME II, SAMM, MMEW, and CAS(ME)3). As seen in Tables 1 to 4, our SKD-TSTSAN demonstrates significant performance improvements in the MER task. Furthermore, the confusion matrices of our SKD-TSTSAN on different datasets are shown in Fig. 5.
As shown in Table 1, we report the performance of SKD-TSTSAN on the CASME II dataset and compare it with other methods. Our SKD-TSTSAN achieves optimal performance on both 3-category and 5-category MER tasks. When performing the 3-category classification, SKD-TSTSAN achieves a superior 99.13% UF1 score and 98.67% UAR, exhibiting a 3.81% UF1 improvement and 3.51% UAR improvement compared to the second-ranked MER method, HTNet (Wang et al.,, 2023). For the 5-category emotion classification, SKD-TSTSAN achieved the highest UF1 of 90.34% and UAR of 89.47%, compared to $\mu$-BERT’s 85.53% UF1 and 83.48% UAR. Furthermore, compared to our baseline TSCNN, SKD-TSTSAN shows a 9.64% UF1 and 8.50% UAR improvement, proving the effectiveness of our submodules.
Table 3 presents the quantitative comparison results for the 3-category MER task on the SAMM dataset. Our SKD-TSTSAN achieves a UF1 of
Table 2: Comparisons with other approaches on SAMM dataset.
Method # Classes UF1 (%) UAR (%)
OFF-ApexNet (Gan et al.,, 2019) 3 54.09 53.92
STSTNet (Liong et al.,, 2019) 3 65.88 68.10
EMR (Liu et al.,, 2019) 3 77.54 71.52
FeatRef (Zhou et al.,, 2022) 3 73.72 71.55
HTNet (Wang et al.,, 2023) 3 81.31 81.24
SKD-TSTSAN (ours) 3 94.29 92.65
@@
94.29% and a UAR of 92.65%, showing an increase of 12.98% and 11.41% compared to the previous leading method HTNet (Wang et al.,, 2023). All these methods are deep learning approaches that utilize optical flow features. The difference is that our SKD-TSTSAN uses the optical flow fields between the three key-frames as the input to the T-stream, while the other methods (Liong et al.,, 2019; Wang et al.,, 2023; Gan et al.,, 2019; Liu et al.,, 2019; Zhou et al.,, 2022) only utilize the optical flow field between the onset frame and the apex frame. Therefore, SKD-TSTSAN can leverage motion information from different temporal regions for temporal modeling, fully considering the temporal information of ME and achieving the best recognition performance.
@@
In addition, the confusion matrices in Fig. 5 report the results of SKD-TSTSAN on multiple datasets. We can see that on the CASME II and SAMM
Figure 5: Confusion matrices for MER with SKD-TSTSAN on CASME II, SAMM, MMEW, and CAS(ME)3 datasets. (a) and (b) denote the 3-class and 5-class classification results on CASME II. (c) and (d) indicate the results of 4-class and 6-class classification on MMEW. (e) shows the results of 3-class classification on SAMM. (f), (g), and (h) represent the 3-class, 4-class, and 7-class classification results on CAS(ME)3, respectively.
datasets, SKD-TSTSAN achieves good recognition results for all categories. On the MMEW dataset, SKD-TSTSAN performs well on the 4-class MER task. On the 6-class MER task, SKD-TSTSAN has poor recognition performance for the “Fear” and “Sadness” samples. One of the main reasons is the small number of samples for the “Fear” and “Sadness” emotion categories. Besides, different emotion categories may correspond to the same AU. For example, both “Fear” and “Surprise” expressions appear in AU5 (“Upper Lid Raiser”), while “Fear” and “Disgust” expressions both involve AU4 (“Brow Lowerer”), as shown in Fig. 6. Therefore, samples labeled as “Fear” are easily misclassified as “Surprise” and “Disgust” categories, which have larger sample sizes. On the CAS(ME)3 dataset, the performance of SKD-TSTSAN is also affected by the imbalanced class distribution. For the 3 and 4 emotion categories, SKD-TSTSAN has poorer recognition performance on “Positive”, which has fewer samples. For the 7 emotion categories, SKD-TSTSAN performs better in the “Disgust”, “Others”, and “Surprise” categories. One of the reasons for the poor performance on the CAS(ME)3
@@ 4.5.1 Ablation Studies
In this section, to illustrate the effect of each component, we conduct ablation studies on the CASME II and MMEW datasets. Initially, a three-stream baseline network is trained using the focal loss ${{L}_{FL}}$. Subsequently, we incrementally introduce submodules into the baseline model, including motion magnification (Mag) modules, efficient channel attention (ECA) modules, temporal shift modules (TSMs), and self-knowledge distillation (SKD).
@@
We introduce a channel attention mechanism by incorporating ECA modules into the L-stream of the network, which improves UF1 and UAR by 1.13% and 0.12% on the CASME II dataset, and 2.48% and 3.89% on the MMEW dataset. Furthermore, to better understand how ECA modules influence the final recognition performance, the gradient-weighted class activation mapping (Grad-CAM) (Selvaraju et al.,, 2017) is utilized for visualization analysis. Fig. 7(a), (b), and (c) represent the samples of three different emotion categories in the CASME II dataset: “Negative (Disgust)”, “Positive (Happiness)”, and “Surprise”. The apex frames are shown in the first row, and the class activation maps (CAMs) of the SKD-TSTSAN with and without the ECA module are shown in the second and third rows. These maps are from the fourth convolutional layer of the S-stream, and the red regions highlight the areas that are crucial for the MER results. According to the Facial Action Coding System (FACS), the “Disgust” expression usually involves AU9 (“Nose Wrinkler”), AU10 (“Upper Lip Raiser”), and AU4 (“Brow Lowerer”) with AU7 (“Lid Tightener”). The “Happiness” expression often involves AU6 (“Cheek Raiser”) and AU12 (“Lip Corner Puller”). The “Surprise” expression is usually found in AU5 (“Upper Lid Raiser”), AU26 (“Jaw Drop”), AU1 (“Inner Brow Raiser”), and AU2 (“Outer Brow Raiser”). As seen in the CAMs in Fig. 7(a), (b), and (c), when ECA modules are used, the facial regions with strong activation are consistent with the AUs defined by FACS. For instance, in Fig. 7(a), the eyebrow region is highlighted; in Fig. 7(b), the cheek and lip corner regions are emphasized; in Fig. 7(c), the eyebrow and chin regions are highlighted. However, when the ECA module is not used, the model tends to focus on regions less relevant to the expressions. For example, the model without ECA modules focuses on the cheek area in Fig. 7(c), which is not relevant to the “Surprise” expression. Therefore, by incorporating ECA modules into the L-stream, we not only suppress redundant information in the L-stream input but also enable the S-stream to focus
@@
Next, the introduction of TSM in the T-stream enables the exchange of temporal information between adjacent frames and also improves the performance on CASME II and MMEW datasets. When both ECA modules and TSMs are used, the performance is better than using only one of them, demonstrating the complementarity of the introduced modules. Furthermore, when the self-knowledge distillation (SKD) is introduced into MER, UF1 and UAR on CASME II are improved by 1.54% and 0.37%, and UF1 and UAR on MMEW are increased by 3.33% and 3.46%, respectively. For an intuitive analysis of the effect of SKD, as shown in Fig. 8, t-distributed stochastic neighbor embedding (t-SNE) is employed to visualize the learned ME features of subjects 1 and 12 in the CASME II dataset. Fig. 8(a) and (b) illustrate the embedding ME features of the network without and with SKD, respectively. It can be easily observed that with the addition of SKD, the clustering within the same category is closer, and the decision boundaries between different categories are clearer, thus the network has better
@@ 4.5.2 More Analysis on the Effect of Transfer Learning
In Fig. 9(a), we provide the MER results on the CASME II and SAMM datasets without and with transfer learning (TL). The two sets of histograms indicate that the introduction of transfer learning effectively improves the MER performance by 1.86% for the UAR of the 5-class MER task on CASME II and 8.89% for the 3-class MER task on SAMM. This is due to the fact that macro-expressions have similar AUs to MEs in expressing emotions. Therefore, we employ transfer learning in all comparative experiments on ME datasets.
@@ 4.5.3 More Analysis on the Choice of Amplification Factors
To evaluate the effect of different amplification factors (AFs) on the performance, we provide the UAR under different AF configurations on the SAMM and MMEW datasets in Fig. 9(b). For the 3-class MER task on the SAMM dataset, our SKD-TSTSAN consistently outperforms the others MER methods ($\leq$81.24UAR) regardless of the AF value, indicating the
@@
4.5.4 More Analysis on the Temporal Domain Distribution
As illustrated in Fig. 10(a), in our comparative experiment configuration, our T-stream inputs are optical flow maps from two different temporal domains, calculated from the onset frame, apex frame, and offset frame. To explore the effect of different temporal distributions of T-stream inputs, as shown in Fig. 10(b) and (c), we first interpolate each ME sequence to 6 and 11 frames using the temporal interpolation model (TIM) (Pfister et al.,, 2011). The optical flow maps are then calculated between adjacent frames, representing the motion information from 5 and 10 different temporal domains, respectively. Subsequently, the optical flow maps from different temporal domains are input into the T-stream, where temporal modeling is performed using TSM. Table 6 provides the performance of our SKD-TSTSAN on the CASME II dataset under different temporal domain distributions. It’s clear that the optimal results are obtained when we use the optical flow between the three key-frames as T-stream input, with an improvement of 2.96% for UF1 and 2.80% for UAR compared to inputting 10 optical flow maps. The
@@ 4.5.4 More Analysis on the Temporal Domain Distribution
main reason is that the onset frame and offset frame represent the start and end of ME, while the apex frame represents the peak intensity of the expression. The optical flow maps between the three key-frames effectively reflect the motion information. After interpolation using TIM, the motion amplitude between adjacent frames is smaller, resulting in less distinct optical flow features and worse results.
@@ 4.5.5 More Analysis on the Performance of Different Classifiers
To verify that deeper classifiers in SKD extract more discriminative features, we compare the performance of each classifier on multiple datasets. As shown in Table 7, the deepest classifier 3/3 achieves the highest UF1 and UAR scores on different datasets. The deeper classifier 2/3 performs better than the shallow classifier 1/3, since the features extracted from the deeper section are more discriminative. Although classifier 2/3 does not perform as well as the deepest classifier 3/3, its performance can also outperform
@@
other MER methods on different datasets, indicating the effectiveness of our SKD-TSTSAN. To visualize the discriminative features extracted by different classifiers, t-SNE is utilized to visualize the embedding ME features of subjects spNO.142 and spNO.176 in the CAS(ME)3 dataset, as shown in Fig. 11. It is obvious that the deeper classifiers have more concentrated clustering of the same category and clearer decision boundaries between different categories, indicating that more discriminative features are extracted.

### v3

@@ 4.3 Evaluation Metrics
For all experiments on the five public datasets, we employ leave-one-subject-out (LOSO) cross-validation. The unweighted accuracy (UAR) and unweighted F1 score (UF1) are used to evaluate the MER results. For convenience, $T{{P}_{i}}$, $F{{P}_{i}}$, and $F{{N}_{i}}$ are used to represent the predicted numbers of true positives, false positives, and false negatives for each class. $UF1$ and $UAR$ are calculated as follows:
@@
where $C$ represents the number of emotion categories in the dataset, and ${{N}_{i}}$ denotes the total number of MEs in the ${{i}_{th}}$ class.
Table 1: Comparisons with other approaches on composite (Full), CASME II, SAMM, and SMIC datasets.
Method Full CASME II SAMM SMIC
UF1 UAR UF1 UAR UF1 UAR UF1 UAR
OFF-ApexNet (Gan et al.,, 2019) 71.96 70.96 87.64 86.81 54.09 53.92 68.17 66.95
STSTNet (Liong et al.,, 2019) 73.53 76.05 83.82 86.86 65.88 68.10 68.01 70.13
CapsuleNet (Van Quang et al.,, 2019) 65.20 65.06 70.68 70.18 62.09 59.89 58.20 58.77
Dual-Inception (Zhou et al.,, 2019) 73.22 72.78 86.21 85.60 58.68 56.63 66.45 67.26
EMR (Liu et al.,, 2019) 78.85 78.24 82.93 82.09 77.54 71.52 74.61 75.30
FeatRef (Zhou et al.,, 2022) 78.38 78.32 89.15 88.73 73.72 71.55 70.11 70.83
HTNet (Wang et al.,, 2023) 86.03 84.75 95.32 95.16 81.31 81.24 80.49 79.05
SKD-TSTSAN (ours) 90.20 90.83 98.58 98.58 87.70 84.00 84.16 85.71
@@ 4.4 Comparative Experiments Results
In this section, we compare the results of our method with other SOTA methods using five public ME datasets (CASME II, SAMM, SMIC, MMEW, and CAS(ME)3). As seen in Tables 1 to 4, our SKD-TSTSAN demonstrates significant performance improvements in the MER task. Furthermore, the confusion matrices of our SKD-TSTSAN on different datasets are shown in Fig. 5.
Table 1 presents the quantitative comparison results for the 3-category MER task on the composite (Full), CASME II, SAMM, and SMIC datasets. Our SKD-TSTSAN achieves state-of-the-art performance on the composite dataset, with a UF1 of 90.20% and a UAR of 90.83%, showing an improvement of 4.17% and 6.08% over the previous leading method HTNet (Wang et al.,, 2023). On the CASME II dataset, SKD-TSTSAN achieves remarkable results, with a UF1 and UAR of 98.58% and 98.58%, respectively. Similarly, on the SAMM and SMIC datasets, our method outperforms the second-best method HTNet, demonstrating its robustness and generalization ability across different datasets. Most methods are deep learning approaches that utilize optical flow features. The key difference is that our SKD-TSTSAN uses the optical flow fields between the three key frames (onset, apex, and offset) as the input to the T-stream, while the other methods (Liong et al.,, 2019; Wang et al.,, 2023; Gan et al.,, 2019; Liu et al.,, 2019; Zhou et al.,, 2022) only utilize the optical flow field between the onset frame and the apex frame. By leveraging motion information from different temporal regions, SKD-TSTSAN fully considers the temporal dynamics of micro-expressions, which achieves superior recognition performance.
As shown in Table 2, we report the results from the single dataset evaluation on CASME II. For the 5-category emotion classification, SKD-TSTSAN achieved the highest UF1 of 90.34% and UAR of 89.47%, compared to $\mu$-BERT’s 85.53% UF1 and 83.48% UAR. Furthermore, compared to our baseline TSCNN, SKD-TSTSAN shows a 9.64% UF1 and 8.50% UAR improvement, proving the effectiveness of our submodules.
@@
Figure 5: Confusion matrices for MER with SKD-TSTSAN on the composite, CASME II, SAMM, SMIC, MMEW, and CAS(ME)3 datasets. (a), (b), (c), and (d) represent the 3-class classification results on the full composite dataset, CASME II, SAMM, and SMIC datasets, respectively. (e) denotes the 5-class classification results on CASME II. (f) and (g) indicate the 4-class and 6-class classification results on MMEW. (h), (i), and (j) show the 3-class, 4-class, and 7-class classification results on CAS(ME)3, respectively.
In addition, the confusion matrices in Fig. 5 report the results of SKD-TSTSAN on multiple datasets. From Fig. 5, it can be observed that SKD-TSTSAN achieves strong performance on the 3-category MER task for the composite dataset. Specifically, the CASME II dataset shows the best recognition results, while the ”Positive” category on the SAMM dataset and the ”Negative” category on the SMIC dataset exhibit relatively weaker performance. On the MMEW dataset, SKD-TSTSAN performs well on the 4-class MER task. On the 6-class MER task, SKD-TSTSAN has poor recognition performance for the “Fear” and “Sadness” samples. One of the main reasons is the small number of samples for the “Fear” and “Sadness” emotion categories. Besides, different emotion categories may correspond to the same AU. For example, both “Fear” and “Surprise” expressions appear in AU5 (“Upper Lid Raiser”), while “Fear” and “Disgust” expressions both involve AU4 (“Brow Lowerer”), as shown in Fig. 6. Therefore, samples labeled as “Fear” are easily misclassified as “Surprise” and “Disgust” categories, which have larger sample sizes. On the CAS(ME)3 dataset, the performance of SKD-TSTSAN is also affected by the imbalanced class distribution. For the 3 and 4 emotion categories, SKD-TSTSAN has poorer recognition performance on “Positive”, which has fewer samples. For the 7 emotion categories, SKD-TSTSAN performs better in the “Disgust”, “Others”, and “Surprise” categories. One of the reasons for the poor performance on the CAS(ME)3
@@ 4.5.1 Ablation Studies
In this section, to illustrate the effect of each component, we conduct ablation studies on the CASME II and MMEW datasets. Initially, a three-stream baseline network is trained using the focal loss ${{L}_{FL}}$. This baseline network is based on the TSCNN architecture but incorporates several modifications to improve performance, including the addition of batch normalization layers after each convolutional layer, replacing the last pooling operation with global average pooling, and introducing a Dropout layer before the final fully connected layer. Subsequently, we incrementally introduce submodules into the baseline model, including motion magnification (Mag) modules, efficient channel attention (ECA) modules, temporal shift modules (TSMs), and self-knowledge distillation (SKD).
@@
We further introduce a channel attention mechanism by incorporating ECA modules into the L-stream of the network. Compared to the baseline, the ECA module improves UF1 and UAR by 2.62% and 3.98% on the CASME II dataset, and by 1.27% and 2.19% on the MMEW dataset. When combined with the motion magnification module, the ECA module further enhances performance, improving UF1 and UAR by 1.13% and 0.12% on the CASME II dataset, and 2.48% and 3.89% on the MMEW dataset. Furthermore, to better understand how ECA modules influence the final recognition performance, the gradient-weighted class activation mapping (Grad-CAM) (Selvaraju et al.,, 2017) is utilized for visualization analysis. Fig. 7(a), (b), and (c) represent the samples of three different emotion categories in the CASME II dataset: “Negative (Disgust)”, “Positive (Happiness)”, and “Surprise”. The apex frames are shown in the first row, and the class activation maps (CAMs) of the SKD-TSTSAN with and without the ECA module are shown in the second and third rows. These maps are from the fourth convolutional layer of the S-stream, and the red regions highlight the areas that are crucial for the MER results. According to the Facial Action Coding System (FACS), the “Disgust” expression usually involves AU9 (“Nose Wrinkler”), AU10 (“Upper Lip Raiser”), and AU4 (“Brow Lowerer”) with AU7 (“Lid Tightener”). The “Happiness” expression often involves AU6 (“Cheek Raiser”) and AU12 (“Lip Corner Puller”). The “Surprise” expression is usually found in AU5 (“Upper Lid Raiser”), AU26 (“Jaw Drop”), AU1 (“Inner Brow Raiser”), and AU2 (“Outer Brow Raiser”). As seen in the CAMs in Fig. 7(a), (b), and (c), when ECA modules are used, the facial regions with strong activation are consistent with the AUs defined by FACS. For instance, in Fig. 7(a), the eyebrow region is highlighted; in Fig. 7(b), the cheek and lip corner regions are emphasized; in Fig. 7(c), the eyebrow and chin regions are highlighted. However, when the ECA module is not used, the model tends to focus on regions less relevant to the expressions. For example, the model without ECA modules focuses on the cheek area in Fig. 7(c), which is not relevant to the “Surprise” expression. Therefore, by incorporating ECA modules into the L-stream, we not only suppress redundant information in the L-stream input but also enable the S-stream to focus
@@
Next, the introduction of TSM in the T-stream enables the exchange of temporal information between adjacent frames. Compared to the baseline, the TSM improves UF1 and UAR by 6.18% and 7.16% on the CASME II dataset, and by 0.54% and 0.23% on the MMEW dataset. These results demonstrate the effectiveness of TSM in capturing temporal dynamics. When combined with the motion magnification module, the TSM further enhances performance on CASME II and MMEW datasets. Additionally, when both ECA modules and TSMs are used, the performance is better than using only one of them, demonstrating the complementarity of the introduced modules. Furthermore, when the self-knowledge distillation (SKD) is introduced into MER, UF1 and UAR on CASME II are improved by 1.54% and 0.37%, and UF1 and UAR on MMEW are increased by 3.33% and 3.46%, respectively. For an intuitive analysis of the effect of SKD, as shown in Fig. 8, t-distributed stochastic neighbor embedding (t-SNE) is employed to visualize the learned ME features of subjects 1 and 12 in the CASME II dataset. Fig. 8(a) and (b) illustrate the embedding ME features of the network without and with SKD, respectively. It can be easily observed that with the addition of SKD, the clustering within the same category is closer, and the decision boundaries between different categories are clearer, thus the network has better
@@ 4.5.2 More Analysis on the Effect of Transfer Learning
In Fig. 9(a), we provide the MER results obtained by training our model on the CASME II and SAMM datasets without and with transfer learning (TL) from the macro-expression dataset CK+. For the CASME II dataset, the introduction of transfer learning improved the UAR of the 5-class MER task by 1.86%. For the SAMM dataset, transfer learning enhanced the UAR of the 3-class MER task by 8.89%. This improvement is attributed to the fact that macro-expressions and micro-expressions share similar AUs in conveying emotions. Consequently, we employ transfer learning in all comparative experiments on ME datasets.
@@ 4.5.3 More Analysis on the Choice of Amplification Factors
To evaluate the effect of different amplification factors (AFs) on the performance, we provide the UAR results obtained by training our model under different AF configurations on the SAMM and MMEW datasets in Fig. 9(b). For the 3-class MER task on the SAMM dataset, our SKD-TSTSAN consistently outperforms the others MER methods ($\leq$81.24UAR) regardless of the AF value, indicating the
@@
4.5.4 More Analysis on the Temporal Domain Distribution
As illustrated in Fig. 10(a), in our comparative experiment configuration, our T-stream inputs are optical flow maps from two different temporal domains, calculated from the onset frame, apex frame, and offset frame. To explore the effect of different temporal distributions of T-stream inputs, as shown in Fig. 10(b) and (c), we first interpolate each ME sequence to 6 and 11 frames using the temporal interpolation model (TIM) (Pfister et al.,, 2011). The optical flow maps are then calculated between adjacent frames, representing the motion information from 5 and 10 different temporal domains, respectively. Subsequently, the optical flow maps from different temporal domains are input into the T-stream, where temporal modeling is performed using TSM. Table 6 provides the performance of our SKD-TSTSAN on the CASME II dataset under different temporal domain distributions. It’s clear that the optimal results are obtained when we use the optical flow between the three key-frames as T-stream input, with an improvement of 2.96% for UF1 and 2.80% for UAR compared to inputting 10 optical flow maps. The main reason is that the onset frame and offset frame represent the start and end of ME, while the apex frame represents the peak intensity of the expression. The optical flow maps between the three key-frames effectively reflect the motion information. After interpolation using TIM, the motion amplitude between adjacent frames is smaller, resulting in less distinct optical flow features and worse results.
@@ 4.5.5 More Analysis on the Performance of Different Classifiers
To verify that deeper classifiers in SKD extract more discriminative features, we compare the performance of each classifier on multiple datasets. As shown in Table 7, the deepest classifier 3/3 achieves the highest UF1 and UAR scores on different datasets. The deeper classifier 2/3 performs better than the shallow classifier 1/3, since the features extracted from the deeper section are more discriminative. Although classifier 2/3 does not perform as well as the deepest classifier 3/3, its performance on different datasets is comparable to or even better than that of other advanced MER methods, indicating the effectiveness of our SKD-TSTSAN. To visualize the discriminative features extracted by different classifiers, t-SNE is utilized to visualize the embedding ME features of subjects spNO.142 and spNO.176 in the CAS(ME)3 dataset, as shown in Fig. 11. It is obvious that the deeper classifiers have more concentrated clustering of the same category and clearer decision boundaries between different categories, indicating that more discriminative features are extracted.